OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -2 x = 0 x = 2 There is no x value, as there is no solution. x can be any value as, there is an infinite number of solutions.

OpenStudy (anonymous):

how does one solve this this

OpenStudy (anonymous):

2x + y = 1 y=1-2x put above value of y in other equation(4x + 2y = -2) 4x+2(1-2x)=-2 4x+2-4x=-2 2=-2 0_0 so no solution

OpenStudy (igreen):

Solve the first equation for \(y\): \(2x + y = 1\) \(y = -2x + 1\) Now we can plug in \(-2x + 1\) into the 2nd equation to get our \(y\) value. \(4x + 2y = -2\) \(4x + 2(-2x + 1) = -2\) \(4x - 4x + 2 = -2\) \(0 + 2 = -2\) The x-value cancels themselves out. There is no solution.

OpenStudy (igreen):

The lines are parallel so the answer is C.

OpenStudy (anonymous):

oh ok but at first i thought it would be D but i get y its not

OpenStudy (igreen):

Well if it had infinitely many solutions both line A and line B would be the same..but they are not..they're parallel, they don't cross.

OpenStudy (anonymous):

oh ok and gee you would make a good teacher i have to say

OpenStudy (igreen):

Lol, I'm only 13..

OpenStudy (anonymous):

lol im 15 haha i fell stupid